# Object-Centric Representation Learning for Enhanced 3D Semantic Scene Graph Prediction

### v1

@@ Appendix A Related Works
Supervised Contrastive Learning. Contrastive learning has recently emerged as a powerful paradigm for self-supervised representation learning. Gutmann et al. [19] laid the theoretical foundation by casting representation learning as density-ratio estimation between target and noise distributions. SimCLR [6] established the practical effectiveness of this paradigm through the NT-Xent objective, while MoCo [20] extended this idea by introducing a momentum encoder and a memory bank to effectively enlarge the set of negatives. Yeh et al. [56] subsequently introduced Decoupled Contrastive Learning (DCL), a variant designed to reduce negative–positive coupling by removing the positive term from the denominator. Contrastive objectives have also been successfully adapted to fully supervised settings: SupCon [23] exploits class labels directly; TSC [28] utilizes deterministic class centres to address long-tailed distributions; and ParCon [7] incorporates parametric centres within the MoCo framework. Motivated by the principles underlying DCL [56] and SupCon [23], we introduce a cross-modal supervised contrastive objective that leverages textual labels to learn a more discriminative object feature space, thereby enhancing predicate prediction accuracy.
@@
3D Point Cloud Understanding. Early research on 3D point-cloud analysis sought affine-invariant representations directly from raw points to support downstream tasks. PointNet [35] introduced a shared multilayer perceptron with max-pool aggregation for global object descriptors, and PointNet++ [36] refined this design through hierarchical farthest-point sampling and local neighbourhood grouping. Due to their computational efficiency, these models remain foundational for various 3D applications, and many 3DSSG pipelines still build upon them. Subsequent work has demonstrated that unsupervised objectives can yield even more transferable point descriptors. PointContrast [53] aligns local fragments from multiple scans of the same scene, while Hou et al. [21] exploit scene-level context by applying contrastive objectives on scene graphs, thereby improving data efficiency on indoor datasets. CrossPoint [1] extends such alignment to the cross-modal setting of point clouds and RGB images, and CLIP2 [60] jointly aligns language, images, and point clouds. The latter, however, employs a fixed prompt vocabulary and does not explicitly address instance-level discrimination within a 3D scene. In this work, we address this limitation by proposing a supervised contrastive formulation that explicitly leverages available image and text annotations. Drawing inspiration from CLIP2 [60], our object encoder fuses visual and spatial cues to produce more discriminative and semantically meaningful instance-level representations.
@@
3D Scene Graph Prediction. Several approaches have attempted to incorporate relational or global context into 3D scene understanding. For instance, the pioneering SGPN [45] combined PointNet [35] with a graph neural network and introduced the widely adopted 3DSSG benchmark. Subsequent graph-based methods propagate features along adjacency edges to refine node embeddings and encode spatial priors, but their effectiveness remains constrained by the quality and discriminative power of initial object descriptors. Within 3D scene-graph generation, SGFN [52] introduced a feature-wise attention mechanism within a GNN framework, substantially outperforming geometry-only baselines relying solely on PointNet embeddings. Similarly, SGGpoint [61] leveraged edge-oriented graph convolutions to incorporate multi-dimensional geometric features explicitly into relationship modeling, while Zhang et al. [63] employed graph auto-encoders to explicitly integrate prior knowledge of object and predicate classes. More recent studies have explored alternative learning paradigms to address limitations inherent to fully supervised training. VL-SAT [48] leverages visual–linguistic auxiliary tasks during training, achieving state-of-the-art performance by enriching object and predicate representations. Koch et al. [26] employ a self-supervised reconstruction objective as a regularizer to enhance the quality of learned latent representations, and 3D-VLAP [47] adopts a weakly supervised approach, aligning 3D data with 2D images and textual labels.
@@
Object-Centric Representation Learning. Object-centric approaches have proliferated in both vision and NLP domains. In 2D scene-graph generation, KERN [5] leverages statistical object co-occurrence patterns to enhance relationship prediction accuracy while regularizing semantic predictions to address uneven distribution issues. Similar to the present framework, KERN relies on high-quality object detections from Faster R-CNN. For visual-relationship detection, ViP [29] implements a phrase-guided message-passing structure that jointly infers subject, predicate, and object elements. In knowledge-graph completion research, TransR [31] maps entities and relations into distinct embedding spaces, extending previous models such as TransE and TransH. The approach proposed in this work builds upon these foundational ideas while placing greater emphasis on learning highly discriminative object features, which subsequently serve as strong priors for predicate estimation.
@@
Open-Vocabulary Settings. Recently, research utilizing VLM models such as OpenSeg [16] and InstructionBLIP [8] for Scene Graph Generation has been actively conducted. Typically, hybrid approaches that utilize both VLM/LLM and GNN have emerged. Open3DSG [27] is a representative example that performs knowledge distillation to lighter PointNet and GCN using well-trained VLMs[16, 8]. Specifically, they extract object features from scene images through OpenSeg and extract relationship features between two objects in images through InstructionBLIP. They perform knowledge distillation so that this information can be reflected in the node/edge features of GNN and the object/relationship PointNet respectively, pretraining the final scene graph model. During inference, they use CLIP for nodes and LLM for estimating relationships between nodes in an open-vocabulary setting. Meanwhile, Gu et al. [18] presented an approach that constructs Scene Graphs using only a combination of VLM/LLM. They primarily utilized LLaVA-7B, an LVLM, to detect objects from RGBD image sequences captured in scenes and set these as nodes in the scene graph. After node detection, they construct an MST (Minimum Spanning Tree) using the IOU of bounding boxes as weights for all 3D object pairs, and finally use LLMs such as GPT-4 to infer edges. By utilizing LVLM/LLM in both node and edge prediction processes, their work became pioneering research in scene graph generation in open-vocabulary settings, demonstrating the potential for expansion to various downstream tasks. Additionally, in 2025, Zhang et al. [62] extended the existing 3D Scene Graph Generation task to propose a VLM/LLM-based model that outperforms existing models, along with a dataset that can construct more practical scene graphs considering interactive objects and functional relationships. Similar other studies [39, 22, 59, 10, 49] are continuing research in directions similar to ConceptGraph [18], showing great potential for development in this field.
@@
Our method departs from prior work in three key aspects. First, whereas most existing approaches primarily focus on predicate prediction, we prioritize learning highly discriminative and semantically robust object-level representations, which subsequently improve predicate inference. Second, in contrast to earlier contrastive pipelines [1, 53, 60], our encoder is explicitly optimized at the instance level, effectively capturing relational semantics critical to the demands of 3DSSG. Lastly, compared to those researches using only VLM/LLM to build scene graph, our study focused on closed-vocabulary settings to effectively verify our insights noted in Section 2, Observation of manuscript. Considering the fundamental differences between GNN-based and VLM/LLM-based methods, we chose a GNN-based approach to clearly and effectively verify our hypothesis: predicate classification errors are strongly correlated with incorrect or uncertain object label predictions. On the other hand, we excluded VLM/LLM from our choices to remove dependency on foundation model performance which will limit our ability to improve object encoder discrimination power.
@@ B.2 Probabilistic Details of hypothesis formulation
Extensibility to open-vocabulary settings. Since we adopt the closed-vocabulary setting of 3DSSG [45], our assumptions yield strong performance on 3D scene graph prediction. By contrast, these assumptions are not directly compatible with open-vocabulary settings. In 3DSSG, predicate semantics primarily encode geometric relations (e.g., ‘hanging on’, ‘standing on’), and prior work correspondingly targets a limited predicate label set. While our results show that this focus is effective under a closed vocabulary, it can overlook novel relations that arise beyond the predefined set. We posit that if object embeddings are made discriminative for additional semantics—such as functionality or color—predicate estimation could be extended toward open-vocabulary settings. In particular, decoupling fine-grained object and predicate categories and aligning the embedding space with this factorization offers a viable path to adapting our framework to open-vocabulary 3D scene graph prediction.
@@
Unlike existing pre-training approaches for 3D point clouds, this work employs supervised contrastive learning in which the positive term is deliberately omitted from the loss denominator. As noted in Appendix § A, prior work on multimodal contrastive learning for 3D objects typically relies on InfoNCE-style losses, similar to those used in SimCLR. The conventional motivation is to maximize mutual information $I(o;\mathbf{z})$ by minimizing the InfoNCE loss, which can be interpreted as the conditional entropy $H(o|\mathbf{z})$ with positive terms included in the denominator. However, as in DCL [56], the positive term in the denominator is deliberately omitted to enhance representational capacity within the self-supervised learning framework. This section provides theoretical evaluation; the precise loss definition is provided in the manuscript. To analyze the impact of the positive term in the denominator, consider a text-modal loss that contains this term, similar to previous works  [23, 1, 60, 34, 7]. $\mathcal{L}^{i}_{\text{text}}$ can be formulated as:
@@
There exists a multiplier $q_{\text{text}}^{i}$ in the gradient $\mathcal{L}_{\text{text}}^{i}$ analogous to the negative-positive coupling (NPC) term proposed in [56].
@@
Proposition 2 demonstrates that retaining the positive term in the denominator reproduces the critical issues identified in [56]: it simplifies the pretext task and generates disproportionately large gradients for instances with close positive and distant negative pairs. In the 3DSSG context, eliminating $q_{\text{text}}^{i}$ provides compelling analytical evidence supporting this hypothesis. Since every $\mathbf{z}_{\text{text}}^{p}$ with the same class label is identical, $q_{\text{text}}^{i}$ can be reformulated as:
@@
Let $c$ denote the class label index and $c_{i}$ the class label index of the anchor. Under this definition, $\mathbf{z}_{\text{text}}^{c_{i}}$ serves as the unique positive for the anchor $\mathbf{z}^{t}_{i}$. The Cross Entropy-style loss is therefore given by Eq.  19, and the corresponding multiplier $q_{\text{CE-like}}^{i}$, following DCL [56], is formulated as:
@@
This formulation clearly exhibits the problems described in DCL [56]. It neither addresses the long-tailed distribution issues highlighted in Eq. 14 nor produces a sufficiently discriminative feature space, making it unsuitable for validating the hypotheses.
@@ Appendix C Implementation Details
Object encoder pre-training. Object encoder pre-training is performed in PyTorch 1.12 (CUDA 11.3) on a single NVIDIA GeForce RTX 3060 Ti GPU. The network is optimized for 100 epochs using Adam, with a global batch size of 512, an initial learning rate of 0.01, and a cosine decay schedule to zero. Training takes roughly five hours to converge, and we retain the checkpoint achieving the highest cumulative validation accuracy summed across top-$K$ accuracies for $K\in\{1,5,10\}$ where classifier is CLIP[37]-based maximum similarity selection. For data preparation, we extract individual object point clouds from the 3DSSG scenes, translate them to the origin, randomly rotate them about the $z$-axis, and uniformly down-sample to 256 points. RGB-D images are cropped according to the VL-SAT pipeline [48]. Using CLIP [37], we select the four most similar images within the same scene that can be re-projected onto the object point cloud. As suitable images are not always available, the visual contrastive loss accommodates multiple images per positive sample. The original 3DSSG train/validation split is preserved throughout pre-training. The object encoder converges after approximately four hours under this setting. To evaluate classification accuracy in isolation, we train a three-layer MLP using Adam with a batch size of 256, a fixed learning rate of $1\times 10^{-4}$, and standard Cross Entropy loss, without additional learning-rate scheduling.
@@ D.1 Metrics
Consistent with the 3DSSG protocol [45], we report top-$K$ recall (R@$K$) for both object and predicate recognition. Triplet scores are computed as the product of the subject, predicate, and object confidences; a triplet is deemed correct only when all three labels match the ground truth, and R@$K$ is then evaluated over these scores. We also adopt the two standard tasks introduced in [54]: Scene Graph Classification (SGCls), which evaluates complete triplets given ground-truth object boxes, and Predicate Classification (PredCls), which assesses predicate predictions under ground-truth object categories and boxes. For both tasks, we follow the evaluation protocol of Zhang et al. [63] and report recall at top-$K$ (R@$K$), counting a prediction as correct if the predicted subject, predicate, and object jointly match at least one ground-truth relation. Collectively, these metrics provide a comprehensive assessment of both object recognition accuracy and relational reasoning performance.

### v2

@@ Appendix A Related Works
Supervised Contrastive Learning. Contrastive learning has recently emerged as a powerful paradigm for self-supervised representation learning. Gutmann et al. [gutmann2010noise] laid the theoretical foundation by casting representation learning as density-ratio estimation between target and noise distributions. SimCLR [chen2020simple] established the practical effectiveness of this paradigm through the NT-Xent objective, while MoCo [he2020momentum] extended this idea by introducing a momentum encoder and a memory bank to effectively enlarge the set of negatives. Yeh et al. [yeh2022decoupled] subsequently introduced Decoupled Contrastive Learning (DCL), a variant designed to reduce negative–positive coupling by removing the positive term from the denominator. Contrastive objectives have also been successfully adapted to fully supervised settings: SupCon [khosla2020supervised] exploits class labels directly; TSC [li2022targeted] utilizes deterministic class centres to address long-tailed distributions; and ParCon [cui2021parametric] incorporates parametric centres within the MoCo framework. Motivated by the principles underlying DCL [yeh2022decoupled] and SupCon [khosla2020supervised], we introduce a cross-modal supervised contrastive objective that leverages textual labels to learn a more discriminative object feature space, thereby enhancing predicate prediction accuracy.
@@
3D Point Cloud Understanding. Early research on 3D point-cloud analysis sought affine-invariant representations directly from raw points to support downstream tasks. PointNet [qi2017pointnet] introduced a shared multilayer perceptron with max-pool aggregation for global object descriptors, and PointNet++ [qi2017pointnet++] refined this design through hierarchical farthest-point sampling and local neighbourhood grouping. Due to their computational efficiency, these models remain foundational for various 3D applications, and many 3DSSG pipelines still build upon them. Subsequent work has demonstrated that unsupervised objectives can yield even more transferable point descriptors. PointContrast [xie2020pointcontrast] aligns local fragments from multiple scans of the same scene, while Hou et al. [hou2021exploring] exploit scene-level context by applying contrastive objectives on scene graphs, thereby improving data efficiency on indoor datasets. CrossPoint [afham2022crosspoint] extends such alignment to the cross-modal setting of point clouds and RGB images, and CLIP2 [zeng2023clip2] jointly aligns language, images, and point clouds. The latter, however, employs a fixed prompt vocabulary and does not explicitly address instance-level discrimination within a 3D scene. In this work, we address this limitation by proposing a supervised contrastive formulation that explicitly leverages available image and text annotations. Drawing inspiration from CLIP2 [zeng2023clip2], our object encoder fuses visual and spatial cues to produce more discriminative and semantically meaningful instance-level representations.
@@
3D Scene Graph Prediction. Several approaches have attempted to incorporate relational or global context into 3D scene understanding. For instance, the pioneering SGPN [wald2020learning] combined PointNet [qi2017pointnet] with a graph neural network and introduced the widely adopted 3DSSG benchmark. Subsequent graph-based methods propagate features along adjacency edges to refine node embeddings and encode spatial priors, but their effectiveness remains constrained by the quality and discriminative power of initial object descriptors. Within 3D scene-graph generation, SGFN [wu2021scenegraphfusion] introduced a feature-wise attention mechanism within a GNN framework, substantially outperforming geometry-only baselines relying solely on PointNet embeddings. Similarly, SGGpoint [zhang2021sggpoint] leveraged edge-oriented graph convolutions to incorporate multi-dimensional geometric features explicitly into relationship modeling, while Zhang et al. [zhang2021knowledge] employed graph auto-encoders to explicitly integrate prior knowledge of object and predicate classes. More recent studies have explored alternative learning paradigms to address limitations inherent to fully supervised training. VL-SAT [wang2023vl] leverages visual–linguistic auxiliary tasks during training, achieving state-of-the-art performance by enriching object and predicate representations. Koch et al. [koch2024sgrec3d] employ a self-supervised reconstruction objective as a regularizer to enhance the quality of learned latent representations, and 3D-VLAP [wang2024weakly] adopts a weakly supervised approach, aligning 3D data with 2D images and textual labels.
@@
Object-Centric Representation Learning. Object-centric approaches have proliferated in both vision and NLP domains. In 2D scene-graph generation, KERN [chen2019knowledge] leverages statistical object co-occurrence patterns to enhance relationship prediction accuracy while regularizing semantic predictions to address uneven distribution issues. Similar to the present framework, KERN relies on high-quality object detections from Faster R-CNN. For visual-relationship detection, ViP [li2017vip] implements a phrase-guided message-passing structure that jointly infers subject, predicate, and object elements. In knowledge-graph completion research, TransR [lin2015learning] maps entities and relations into distinct embedding spaces, extending previous models such as TransE and TransH. The approach proposed in this work builds upon these foundational ideas while placing greater emphasis on learning highly discriminative object features, which subsequently serve as strong priors for predicate estimation.
@@
Open-Vocabulary Settings. Recently, research utilizing VLM models such as OpenSeg [ghiasi2022scaling] and InstructionBLIP [dai2023instructblip] for Scene Graph Generation has been actively conducted. Typically, hybrid approaches that utilize both VLM/LLM and GNN have emerged. Open3DSG [koch2024open3dsg] is a representative example that performs knowledge distillation to lighter PointNet and GCN using well-trained VLMs[ghiasi2022scaling, dai2023instructblip]. Specifically, they extract object features from scene images through OpenSeg and extract relationship features between two objects in images through InstructionBLIP. They perform knowledge distillation so that this information can be reflected in the node/edge features of GNN and the object/relationship PointNet respectively, pretraining the final scene graph model. During inference, they use CLIP for nodes and LLM for estimating relationships between nodes in an open-vocabulary setting. Meanwhile, Gu et al. [gu2024conceptgraphs] presented an approach that constructs Scene Graphs using only a combination of VLM/LLM. They primarily utilized LLaVA-7B, an LVLM, to detect objects from RGBD image sequences captured in scenes and set these as nodes in the scene graph. After node detection, they construct an MST (Minimum Spanning Tree) using the IOU of bounding boxes as weights for all 3D object pairs, and finally use LLMs such as GPT-4 to infer edges. By utilizing LVLM/LLM in both node and edge prediction processes, their work became pioneering research in scene graph generation in open-vocabulary settings, demonstrating the potential for expansion to various downstream tasks. Additionally, in 2025, Zhang et al. [zhang2025open] extended the existing 3D Scene Graph Generation task to propose a VLM/LLM-based model that outperforms existing models, along with a dataset that can construct more practical scene graphs considering interactive objects and functional relationships. Similar other studies [rotondi2025fungraph, kassab2024bare, zemskova20243dgraphllm, devarakonda2024orionnav, werby2024hierarchical] are continuing research in directions similar to ConceptGraph [gu2024conceptgraphs], showing great potential for development in this field.
@@
Our method departs from prior work in three key aspects. First, whereas most existing approaches primarily focus on predicate prediction, we prioritize learning highly discriminative and semantically robust object-level representations, which subsequently improve predicate inference. Second, in contrast to earlier contrastive pipelines [afham2022crosspoint, xie2020pointcontrast, zeng2023clip2], our encoder is explicitly optimized at the instance level, effectively capturing relational semantics critical to the demands of 3DSSG. Lastly, compared to those researches using only VLM/LLM to build scene graph, our study focused on closed-vocabulary settings to effectively verify our insights noted in Section 2, Observation of manuscript. Considering the fundamental differences between GNN-based and VLM/LLM-based methods, we chose a GNN-based approach to clearly and effectively verify our hypothesis: predicate classification errors are strongly correlated with incorrect or uncertain object label predictions. On the other hand, we excluded VLM/LLM from our choices to remove dependency on foundation model performance which will limit our ability to improve object encoder discrimination power.
@@ B.2 Probabilistic Details of hypothesis formulation
Extensibility to open-vocabulary settings. Since we adopt the closed-vocabulary setting of 3DSSG [wald2020learning], our assumptions yield strong performance on 3D scene graph prediction. By contrast, these assumptions are not directly compatible with open-vocabulary settings. In 3DSSG, predicate semantics primarily encode geometric relations (e.g., ‘hanging on’, ‘standing on’), and prior work correspondingly targets a limited predicate label set. While our results show that this focus is effective under a closed vocabulary, it can overlook novel relations that arise beyond the predefined set. We posit that if object embeddings are made discriminative for additional semantics—such as functionality or color—predicate estimation could be extended toward open-vocabulary settings. In particular, decoupling fine-grained object and predicate categories and aligning the embedding space with this factorization offers a viable path to adapting our framework to open-vocabulary 3D scene graph prediction.
@@
Unlike existing pre-training approaches for 3D point clouds, this work employs supervised contrastive learning in which the positive term is deliberately omitted from the loss denominator. As noted in Appendix § A, prior work on multimodal contrastive learning for 3D objects typically relies on InfoNCE-style losses, similar to those used in SimCLR. The conventional motivation is to maximize mutual information $I(o;\mathbf{z})$ by minimizing the InfoNCE loss, which can be interpreted as the conditional entropy $H(o|\mathbf{z})$ with positive terms included in the denominator. However, as in DCL [yeh2022decoupled], the positive term in the denominator is deliberately omitted to enhance representational capacity within the self-supervised learning framework. This section provides theoretical evaluation; the precise loss definition is provided in the manuscript. To analyze the impact of the positive term in the denominator, consider a text-modal loss that contains this term, similar to previous works  [khosla2020supervised, afham2022crosspoint, zeng2023clip2, oord2018representation, cui2021parametric]. $\mathcal{L}^{i}_{\text{text}}$ can be formulated as:
@@
There exists a multiplier $q_{\text{text}}^{i}$ in the gradient $\mathcal{L}_{\text{text}}^{i}$ analogous to the negative-positive coupling (NPC) term proposed in [yeh2022decoupled].
@@
Proposition 2 demonstrates that retaining the positive term in the denominator reproduces the critical issues identified in [yeh2022decoupled]: it simplifies the pretext task and generates disproportionately large gradients for instances with close positive and distant negative pairs. In the 3DSSG context, eliminating $q_{\text{text}}^{i}$ provides compelling analytical evidence supporting this hypothesis. Since every $\mathbf{z}_{\text{text}}^{p}$ with the same class label is identical, $q_{\text{text}}^{i}$ can be reformulated as:
@@
Let $c$ denote the class label index and $c_{i}$ the class label index of the anchor. Under this definition, $\mathbf{z}_{\text{text}}^{c_{i}}$ serves as the unique positive for the anchor $\mathbf{z}^{t}_{i}$. The Cross Entropy-style loss is therefore given by Eq.  19, and the corresponding multiplier $q_{\text{CE-like}}^{i}$, following DCL [yeh2022decoupled], is formulated as:
@@
This formulation clearly exhibits the problems described in DCL [yeh2022decoupled]. It neither addresses the long-tailed distribution issues highlighted in Eq. 14 nor produces a sufficiently discriminative feature space, making it unsuitable for validating the hypotheses.
@@ Appendix C Implementation Details
Object encoder pre-training. Object encoder pre-training is performed in PyTorch 1.12 (CUDA 11.3) on a single NVIDIA GeForce RTX 3060 Ti GPU. The network is optimized for 100 epochs using Adam, with a global batch size of 512, an initial learning rate of 0.01, and a cosine decay schedule to zero. Training takes roughly five hours to converge, and we retain the checkpoint achieving the highest cumulative validation accuracy summed across top-$K$ accuracies for $K\in\{1,5,10\}$ where classifier is CLIP[radford2021learning]-based maximum similarity selection. For data preparation, we extract individual object point clouds from the 3DSSG scenes, translate them to the origin, randomly rotate them about the $z$-axis, and uniformly down-sample to 256 points. RGB-D images are cropped according to the VL-SAT pipeline [wang2023vl]. Using CLIP [radford2021learning], we select the four most similar images within the same scene that can be re-projected onto the object point cloud. As suitable images are not always available, the visual contrastive loss accommodates multiple images per positive sample. The original 3DSSG train/validation split is preserved throughout pre-training. The object encoder converges after approximately four hours under this setting. To evaluate classification accuracy in isolation, we train a three-layer MLP using Adam with a batch size of 256, a fixed learning rate of $1\times 10^{-4}$, and standard Cross Entropy loss, without additional learning-rate scheduling.
@@ D.1 Metrics
Consistent with the 3DSSG protocol [wald2020learning], we report top-$K$ recall (R@$K$) for both object and predicate recognition. Triplet scores are computed as the product of the subject, predicate, and object confidences; a triplet is deemed correct only when all three labels match the ground truth, and R@$K$ is then evaluated over these scores. We also adopt the two standard tasks introduced in [xu2017scene]: Scene Graph Classification (SGCls), which evaluates complete triplets given ground-truth object boxes, and Predicate Classification (PredCls), which assesses predicate predictions under ground-truth object categories and boxes. For both tasks, we follow the evaluation protocol of Zhang et al. [zhang2021knowledge] and report recall at top-$K$ (R@$K$), counting a prediction as correct if the predicted subject, predicate, and object jointly match at least one ground-truth relation. Collectively, these metrics provide a comprehensive assessment of both object recognition accuracy and relational reasoning performance.